# Recursive Inference Scaling: A Winning Path to Scalable Inference in Language and Multimodal Systems

### v2

@@ 4 Stochastic Recursive Inference Scaling
Figure 4: Performance of RINS (A2B, A3B, A4B) vs. RAO (A2, A3, A4) with increasing compute on 600M-parameter models. All models have the same size. The performance advantage of RINS grows with the computational budget. The long-sequence baseline is not shown since it underperforms other methods in Figure 2.
Next, we investigate the effect of stochastically dropping blocks during training, inspired by the regularization technique of stochastic depth (Huang et al., 2016). Our primary goal is to determine whether this approach can further enhance the performance of RINS while simultaneously offering the flexibility of reverting to non-recursive inference without significant degradation in model quality.
To recall, RINS has the signature ArB for some $r>1$. To implement stochastic RINS, we introduce a skip probability $p_{s}\in[0,1)$ and sample during training the number of recursion rounds in each step to be $1+\eta$, where $\eta$ is a binomial random variable with probability of success $1-p_{s}$ and number of trials $r-1$. Thus, block A is always executed at least once. During inference, we are free to choose how to scale inference compute by setting the value of $r\geq 1$. See the detailed pseudocode in Figure 3.
Here, we train bigger models with 1 billion parameters. We use an embedding dimension 2,048 and MLP dimension 8,192. All models have 18 decoder blocks. We train for 500K steps and compare RINS with signature A3B against the non-recursive baseline.
Figure 5 summarizes the advantages of stochastic RINS. Notably, we observe that as $p_{s}>0$ increases, stochastic RINS mitigates the performance degradation incurred when scaling is not applied at inference time, while still offering big gains when inference time is scaled. Not surprisingly, though, scaling inference time is less effective when $p_{s}$ increases, suggesting a tradeoff between flexibility at inference time and the magnitude of potential gains from scaling. As shown in Figure 6, similar conclusions hold even in the asymptotic (infinite-compute) regime, under the assumption that loss follows a power law relation (Kaplan et al., 2020).
Figure 5: Performance of stochastic RINS (A3B) with varying inference costs for 1B parameter LMs. The $x$-axis represents the training compute cost. The legend indicates the inference cost of each stochastic RINS configuration relative to the baseline; e.g. $1.5x$ denotes 50% increase in inference cost. For $p_{s}=0$, RINS@1x is significantly worse, with perplexity scores $>3$. As expected, RINS converges in performance to the baseline as $p_{s}\to 1$.
Figure 6: Asymptotic performance of 1B-parameter LMs, evaluated in C4 (left) and SlimPajama (right). Dotted line is for baseline. We observe a tradeoff between inference flexibility (gap between 1x & 2x) and magnitude of gain from inference scaling (2x result).
5 Other Modalities
5.1 Vision
Setup. We train an encoder-only vision transformer ViT-B/16 (Dosovitskiy et al., 2021) on ImageNet-ILSRCV2012 (Deng et al., 2009). The non-recursive baseline model is trained for either 300 or 1,000 epochs using a batch size 1,024, while recursive models are trained on fewer epochs to match the same total training compute FLOPs. We apply MixUp (probability 0.2) during training (Zhang et al., 2018) and use learned position embedding. The optimizer is Adam where we tune the learning rate for each architecture in the set $\mathrm{lr}\in\{10^{-3},7\times 10^{-4},3\times 10^{-4},10^{-4}\}$ with weight decay $\mathrm{wd}=\mathrm{lr}/10$, on a small validation split. We use a cosine learning rate schedule with 10K warm-up steps. Images are resized to $224\times 224$ and $16\times 16$ patch sizes are used. The full training configuration is in Appendix A.
Architecture Val ReaL v2 Avg
300 epochs
(ABB)2 75.2 81.4 62.7 73.1
A@336 75.7 81.0 62.3 73.0
AAB 75.2 80.5 61.4 72.4
ABBC 75.1 80.2 61.3 72.2
A@224 74.9 80.1 61.3 72.1
1,000 epochs
A@336 77.6 82.7 64.6 75.0
ABBC 76.3 81.6 63.2 73.7
AABC 76.0 81.4 62.4 73.2
AA 75.8 81.4 62.4 73.3
AAA 75.7 80.8 62.0 72.8
5.2 Contrastive Mutlimodal Systems
Setup.
Finally, we also study the impact of RINS in vision-language pretraining, motivated by the fact that such models also process natural language. We pretrain SigLIP-B/16 models (Zhai et al., 2023), which are contrastive models trained using the sigmoid loss on English-language image–text pairs. We follow (Zhai et al., 2023) in most aspects. Images are resized to $256\times 256$ and we use $16\times 16$ patch sizes. Texts, on the other hand, are tokenized using C4 tokenizer (Raffel et al., 2020) with a vocabulary size of 32K, and we keep a maximum of 64 tokens. The optimizer is Adam with learning rate $10^{-3}$ and weight decay $10^{-4}$, using an inverse square root decay schedule with 5K warm-up and cool-down steps. For the baseline, we use SigLIP-B/16 pretrained on 10B training examples. Again, recursive models are trained on fewer examples to match the total training compute cost. Due to the amount of compute involved in these experiments, we only compare the non-recursive baseline (with signature AB) against RAO (with signature ABAB) and RINS (signature A2B) with degree 1.
Metric Architecture
AB Long-Seq ABAB AAB
Zero-shot classification
ImageNet 73.4 73.0 72.7 74.1
CIAFR100 68.9 63.6 65.3 72.2
Pet 90.4 90.0 90.7 90.0
Retrieval
COCO img2txt@1 62.7 62.0 61.1 62.3
COCO img2txt@5 84.8 84.1 82.9 84.1
COCO img2txt@10 90.7 90.4 89.5 90.2
COCO txt2img@1 44.6 44.2 43.2 45.1
COCO txt2img@5 69.6 69.4 68.4 70.0
COCO txt2img@10 78.8 78.7 77.5 79.0
Flickr img2txt@1 89.6 87.2 87.9 88.9
Flickr img2txt@5 98.0 97.8 97.5 98.5
Flickr img2txt@10 99.1 98.9 98.9 99.3
Flickr txt2img@1 75.1 73.7 73.2 74.3
Flickr txt2img@5 92.3 92.1 91.5 92.4
Flickr txt2img@10 95.6 95.6 95.6 95.8
Table 3: Performance of SigLIP-B/16 on various datasets. Results are shown for ImageNet (Deng et al., 2009), CIFAR100 (Krizhevsky, 2009), Pet (Parkhi et al., 2012), COCO (Chen et al., 2015), and Flickr (Young et al., 2014). All models are identical in size to SigLIP-B/16 and have the same training compute FLOPs. Long-Seq is SigLIP-B/16 trained on higher resolution of 280 and text length 80 (25% increase in sequence length $\rightarrow$ 50% increase in inference cost compared to baseline, similar to AAB). Using Wilcoxon signed rank test (Wilcoxon, 1992), we obtain $p=0.003$ so the evidence in favor of RINS is statistically significant at the 99% confidence level.
Overtraining. Next, we demonstrate that RINS can substantially advance state-of-the-art results in multimodal systems for a given model size. We train a recursive variant of SigLIP-B/16, denoted SigLIP-RINS-B/16, using the signature A3B. In light of the findings in Section 6, we increase the number of recursions here given the increase in compute.
We adhere to the training protocol outlined above, with the exception that SigLIP-RINS-B/16 is now trained on 40B examples, matching the training data scale of the widely-used, publicly available SigLIP-B/16 checkpoint. Note that both models have an identical size. Moreover, following Pouget et al. (2024), we utilize a training mixture comprising both English and multilingual data to enhance cultural diversity, and report the cultural metrics recommended in Pouget et al. (2024) as well as multilinguality evaluations using XM3600 dataset (Thapliyal et al., 2022). So, to ensure an appropriate comparison of results, we re-train SigLIP-B/16 on 40B examples from the same data mixture. The primary datasets for cultural diversity evaluation are Dollar Street (Rojas et al., 2022), GeoDE (Ramaswamy et al., 2024), and Google Landmarks Dataset v2 (GLDv2) (Weyand et al., 2020). We use the Gemma tokenizer (Mesnard et al., 2024).
Metric SigLIP-B/16 SigLIP-RINS-B/16
(370M params) (370M params)
$p_{s}=0$ $\frac{1}{4}$ $\frac{1}{2}$
Zero-shot classification
ImageNet 77.3 79.0 79.6 79.2
CIAFR100 70.3 78.5 80.7 81.7
Pet 92.6 94.8 94.4 92.7
Cultural Diversity
GeoLoc:Dollar Street 17.6 19.7 19.2 19.3
GeoLoc:GeoDE-Country 22.5 23.3 24.7 23.9
GeoLoc:GeoDE-Region 36.1 37.7 39.7 38.7
Dollar Street 51.5 52.9 53.1 53.1
GeoDE 93.1 93.7 94.3 94.2
GLDv2 51.6 53.7 52.4 52.5
Multilinguality
XM3600 img2txt@1 48.4 53.5 52.6 51.8
XM3600 img2txt@5 68.4 72.6 72.0 70.9
XM3600 img2txt@10 74.4 78.1 77.5 76.5
XM3600 txt2img@1 39.5 43.0 43.1 41.9
XM3600 txt2img@5 59.5 63.0 63.1 61.6
XM3600 txt2img@10 66.1 69.3 69.3 68.0
Retrieval
COCO img2txt@1 67.6 69.4 70.0 69.5
COCO img2txt@5 87.2 88.6 88.9 88.3
COCO img2txt@10 92.6 93.2 93.5 93.2
COCO txt2img@1 50.3 51.5 52.4 52.0
COCO txt2img@5 74.7 75.2 76.0 75.7
COCO txt2img@10 82.6 83.0 83.5 83.2
Flickr img2txt@1 91.9 92.9 93.5 92.4
Flickr img2txt@5 99.3 98.7 99.5 98.7
Flickr img2txt@10 99.6 99.5 99.7 99.3
Flickr txt2img@1 80.1 80.7 81.4 80.5
Flickr txt2img@5 94.6 94.3 95.4 95.0
Flickr txt2img@10 97.1 97.0 97.4 97.0
Table 4: Performance of multilingual SigLIP models on various datasets under an overtraining regime, where training compute is not a constraint. All models are identical in size to SigLIP-B/16. As shown in the rightmost columns, stochastic RINS ($p_{s}>0$) outperforms the other architectures. Detailed results on multilinguality are provided in Appendix B.
6 Further Analysis
6.1 Data scaling laws
Next, we investigate the influence of Recursive INference Scaling (RINS) on the data scaling behavior of language models. Specifically, we fit a power law of the form $\varepsilon(x)=\beta x^{-c}+\varepsilon_{\infty}$, to the log-perplexity loss $\varepsilon(x)$ as a function of the training FLOPs $x$. This allows us to analyze the impact of RINS on both the scaling exponent $c$ and the asymptotic performance limit $\varepsilon_{\infty}$, revealing whether the performance gains in RINS stem from an improved scaling exponent, a lower asymptotic limit, or a combination of both. We use the 600M-parameter language models in Figure 4 whose signature is ArB, for $r\in\{1,2,3,4\}$, where $r=1$ corresponds to the non-recursive baseline. As shown in Figure 7, RINS improves both the scaling exponent and the asymptotic limit. The improvement in the asymptotic limit provides further evidence that the performance gap in favor of RINS is not closed by overtraining non-recursive models.
Figure 7: top: The scaling exponent $c$ in models with signature ArB, where $r=1$ corresponds to the non-recursive baseline. RINS ($r>1$) improves scaling exponents. bottom: Asymptotic log-perplexity $\varepsilon_{\infty}$ is plotted against $r$. RINS improves $\varepsilon_{\infty}$, indicating that overtraining the baseline cannot match its performance gains.
6.2 Optimal Number of Recursion Rounds
We observe from the scaling parameters in the previous section that if the performance of RINS is modeled as a power law $f_{r}(x)=\beta_{r}x^{-c_{r}}+\varepsilon_{r}$, then $c_{r}$ increases with $r$ while $\varepsilon_{r}$ decreases. Furthermore, the coefficient $\beta_{r}$ also increases with $r$. This implies that while scaling inference by increasing $r$ might initially exhibit a higher loss due to the larger $\beta_{r}$, its faster convergence ($c_{r}$) and lower asymptotic limit ($\varepsilon_{r}$) will eventually lead to superior performance. In other words, using the higher recursion level is advantageous eventually, which is consistent with the experimental results.
Figure 8: Optimal number of recursions $r$ in RINS is plotted vs. training compute for 300M-, 600M-, and 1B-parameter LMs.
To observe this more explicitly, we train language models with four signatures ArB: $r\in\{1,2,3,4\}$. Then, we plot the optimal value of $r$ against training compute. As shown from the results in Figure 8, the optimal value of $r$ monotonically increases with training compute, in agreement with earlier results. Also, smaller models benefit more from RINS.
7 Discussion and Related Works
The premise that extending inference time enhances the quality of language model outputs finds support in cognitive science. Human deliberation, often associated with System 2 thinking, is linked to improved decision-making capabilities (Lawson et al., 2020). Mirroring this, numerous computational strategies have been developed to scale inference within Large Language Models (LLMs). Prompt-based methods like Chain-of-Thought (CoT) (Wei et al., 2024) and critique-then-generate approaches (Ankner et al., 2024) are prominent examples. Other iterative refinement methods like ReAct (Yao et al., 2023), Self-Refine (Madaan et al., 2024), and Reflexion, which incorporate feedback and reflection, also improve inference quality. Even simpler strategies, capitalizing on the stochastic nature of LLM decoding that generate multiple responses and select among them, such as self-consistency (Wang et al., 2023), have shown efficacy that scales well across multiple orders of magnitude of inference compute (Brown et al., 2024).
However, the assumption of monotonic improvement with increased inference calls is not guaranteed to hold. As argued by Chen et al. (Chen et al., 2024), repeated sampling may lead to convergence towards the most probable, but not necessarily the optimal, solution. This is particularly pertinent for challenging instances where the probability of a correct answer is below chance (i.e., $<0.5$).
In this work, we introduce a complementary approach, called Recursive INference Scaling (RINS), which can be integrated with other techniques.. RINS identifies a specific form of model recursion as an effective inference scaling strategy, demonstrating significant performance gains over various other recursive architectures, including the Repeat All Over (RAO) strategy identified by Liu et al. (2024) as state-of-the-art for mobile LLMs.
Recursion is a form of parameter sharing, a technique that has been explored for a while. Illustrative examples include Universal Transformers (Dehghani et al., 2019) and ALBERT (Lan et al., 2020). Despite their innovative design, their scaling exponents were smaller than in the vanilla transformer (Tay et al., 2022a) so they failed to subsume it (Tay et al., 2022b). RINS, by contrast, improves both the scaling exponents and asymptotic performance limits.
We hypothesize that the efficacy of RINS in language tasks stems from its exploitation of language’s self-similarity (Alabdulmohsin et al., 2024a), since the recursive block in RINS is a form of scale-invariant decoding (see Appendix D for further discussion). One evidence for this is the fact that RINS does not help in the pure vision domain, but it can still improve performance in multimodal systems that also process language, such as contrastive models. We obtain significant performance gains in SigLIP, for example.
Impact Statement
This paper presents work whose goal is to advance the field of artificial intelligence, specifically in the area of language modeling and multimodal systems. There are many potential societal consequences of advancing AI, both positive, such as improved accessibility to high-quality healthcare and education, and negative, if such systems are misused. Our intention with RINS is to contribute to the ongoing effort to improve the performance and capabilities of these models. While we recognize the general ethical considerations that accompany the advancement of language models and AI, we do not feel that this particular work raises unique or unaddressed ethical concerns beyond the established considerations within the field.
Acknowledgement
We thank Michael Tschannen and Jeremiah Harmsen for their insightful reviews and suggestions, and Google DeepMind at large for providing a supportive research environment.
Appendix A Training Configurations
Appendix B Multilinguality Detailed Results
@@
Table 5: Per-language retrieval evaluations using the Crossmodal-3600 dataset (Thapliyal et al., 2022).
@@ Appendix C Zero-shot Evaluation on Common Sense Reasoning Tasks
Hi Ibrahim, I’ve read your code and here’s a detailed description of the evaluation setup for a research paper:
Evaluation Setup for Zero-Shot Common Sense Reasoning
This section describes the evaluation setup for assessing the zero-shot common sense reasoning capabilities of the language model.
@@
Each model is evaluated on four benchmarks:
  • WinoGrande: A multiple-choice dataset focusing on commonsense knowledge (Sakaguchi et al., 2019).
  • OpenBookQA: A multiple-choice question answering dataset requiring knowledge from an open book of elementary level science facts (Mihaylov et al., 2018).
  • BoolQ: A yes/no question answering dataset requiring the model to determine the truthfulness of a given statement based on a provided passage (Clark et al., 2019).
  • PIQA: A multiple-choice benchmark focusing on physics-related reasoning, such as how we interact with objects in daily life (Bisk et al., 2020).
  • SIQA: A multiple-choice dataset focusing on social common sense reasoning (Sap et al., 2019).
@@
We do not apply any prompting, except in PIQA where we formulate each sentence in the form The goal is: {goal} The solution is: {sol}. We do this in PIQA because, otherwise, the sentences can be difficult to understand, even for humans. In one example, for instance, the goal is: ”Deep clean coffee grinder.” and the two possible solutions are: ”Scrape with rice.” and ”Scrape with flour.” Concatenating directly would result in sentences like ”Deep clean coffee grinder. Scrape with rice.” whose are unclear.
@@ Appendix D Recursion and Self-Similarity
D.1 Architecture Sweep
@@
D.2 Training Configuration
D.2.1 Language Modeling
@@
D.2.2 Vision: Image Classification
@@
D.2.3 Language-Image Pretraining

### v3

@@ Appendix A Recursive Inference Scaling (RINS) Pseudocode
Appendix B Retrieval and Multilinguality Detailed Results
The following tables provide the full retrieval and multilinguality results for SigLIP-RINS-B/16.
@@ Appendix B Retrieval and Multilinguality Detailed Results
Table 4: Per-language retrieval evaluations using the Crossmodal-3600 dataset Thapliyal et al., (2022).
Metric SigLIP-B/16 SigLIP-RINS-B/16
(370M params) (370M params)
$p_{s}=0$ $\frac{1}{4}$ $\frac{1}{2}$
Multilinguality
XM3600 img2txt@1 48.4 53.5 52.6 51.8
XM3600 img2txt@5 68.4 72.6 72.0 70.9
XM3600 img2txt@10 74.4 78.1 77.5 76.5
XM3600 txt2img@1 39.5 43.0 43.1 41.9
XM3600 txt2img@5 59.5 63.0 63.1 61.6
XM3600 txt2img@10 66.1 69.3 69.3 68.0
Retrieval
COCO img2txt@1 67.6 69.4 70.0 69.5
COCO img2txt@5 87.2 88.6 88.9 88.3
COCO img2txt@10 92.6 93.2 93.5 93.2
COCO txt2img@1 50.3 51.5 52.4 52.0
COCO txt2img@5 74.7 75.2 76.0 75.7
COCO txt2img@10 82.6 83.0 83.5 83.2
Flickr img2txt@1 91.9 92.9 93.5 92.4
Flickr img2txt@5 99.3 98.7 99.5 98.7
Flickr img2txt@10 99.6 99.5 99.7 99.3
Flickr txt2img@1 80.1 80.7 81.4 80.5
Flickr txt2img@5 94.6 94.3 95.4 95.0
Flickr txt2img@10 97.1 97.0 97.4 97.0
Table 5: Performance of multilingual SigLIP models on various datasets under an overtraining regime. All models are identical in size to SigLIP-B/16. As shown in the rightmost columns, stochastic RINS ($p_{s}>0$) outperforms the other models. Full retrieval & multilinguality results are in Appendix B.
@@ Appendix C Zero-shot Evaluation on Common Sense Reasoning Tasks
Each model is evaluated on six benchmarks:
  • OpenBookQA: A multiple-choice question answering dataset requiring knowledge from an open book of elementary level science facts (Mihaylov et al.,, 2018).
  • BoolQ: A yes/no question answering dataset requiring the model to determine the truthfulness of a given statement based on a provided passage (Clark et al.,, 2019).
  • PIQA: A multiple-choice benchmark focusing on physics-related reasoning, such as how we interact with objects in daily life (Bisk et al.,, 2020).
  • SIQA: A multiple-choice dataset focusing on social common sense reasoning (Sap et al.,, 2019).
  • HellaSwag: A multiple-choice dataset focusing on sentence completion (Zellers et al.,, 2019).
  • CommonSenseQA: A multiple-choice question answering dataset with distractors (Talmor et al.,, 2019).
@@
We do not apply any prompting, except in BoolQ and PIQA. In BoolQ, we use the prompt template:
    <data[’passage’]> Based on this, the answer to the question:
    <data[’question’]>, is: ...,
which we have found to improve performance significantly.
In PIQA, we formulate each sentence in the form:
The goal is: {goal} The solution is: {sol}.
We do this in PIQA because, otherwise, the sentences can be difficult to understand, even for humans. In one example, for instance, the goal is: "Deep clean coffee grinder." and the two possible solutions are: "Scrape with rice." and "Scrape with flour." Concatenating directly would result in sentences like "Deep clean coffee grinder. Scrape with rice." whose are unclear.
@@ Appendix D Recursion and Self-Similarity
Appendix E Configuration
E.1 Architecture Sweep
@@
E.2 Training Configuration
E.2.1 Language Modeling
@@
E.2.2 Vision: Image Classification
@@
E.2.3 Language-Image Pretraining
@@
Appendix F C4 Evaluation Results with Stochastic RINS
Figure 9: Performance of stochastic RINS (A3B) with varying inference costs for 1B parameter LMs on C4, similar to Figure 4 reported on SlimPajama. The $x$-axis represents the training compute cost. The legend indicates the inference cost of each stochastic RINS configuration relative to the baseline; e.g. $1.5x$ denotes 50% increase in inference cost. For $p_{s}=0$, RINS@1x is significantly worse, with perplexity scores $>3$. As expected, RINS converges in performance to the baseline as $p_{s}\to 1$.
Appendix G Vision
Setup. We train an encoder-only vision transformer ViT-B/16 Dosovitskiy et al., (2021) on ImageNet-ILSRCV2012 Deng et al., (2009). The non-recursive baseline model is trained for either 300 or 1,000 epochs using a batch size 1,024, while recursive models are trained on fewer epochs to match the same total training compute FLOPs. We apply MixUp (probability 0.2) during training Zhang et al., (2018) and use learned position embedding. The optimizer is Adam where we tune the learning rate for each architecture in the set $\mathrm{lr}\in\{10^{-3},7\times 10^{-4},3\times 10^{-4},10^{-4}\}$ with weight decay $\mathrm{wd}=\mathrm{lr}/10$, on a small validation split. We use a cosine learning rate schedule with 10K warm-up steps. Images are resized to $224\times 224$ and $16\times 16$ patch sizes are used. The full training configuration is in Appendix E.
Architecture Val ReaL v2 Avg
300 epochs
(ABB)2 75.2 81.4 62.7 73.1
A@336 75.7 81.0 62.3 73.0
AAB 75.2 80.5 61.4 72.4
ABBC 75.1 80.2 61.3 72.2
A@224 74.9 80.1 61.3 72.1
1,000 epochs
A@336 77.6 82.7 64.6 75.0
ABBC 76.3 81.6 63.2 73.7
AABC 76.0 81.4 62.4 73.2
AA 75.8 81.4 62.4 73.3
AAA 75.7 80.8 62.0 72.8